# Injectives obstruct Fourier-Mukai functors

### v1

@@ 3.2. The $\mathrm{Inj}$-equivariant characteristic morphism
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives. Then the inclusion $\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is dg.
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives. Then the functor $\iota^{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$ is a quasi-equivalence.
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}$ is an isomorphism.
@@
By Section 3.1 above we can interpret the equivariant characteristic morphism as the pullback along the inclusion of the full subcategory $\iota:\mathrm{Inj}\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ in the homotopy category of $\mathrm{dg}$-categories. In particular we need to pick a fibrant replacement of $\mathrm{Inj}$. By [Tab07, Lemme 2.13] pretriangulated replacements are an incarnation of a fibrant replacement. So we may choose the canonical pretriangulated hull given by $\mathrm{dg}$-modules $\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$. By Proposition 3.4 $\iota$ induces an equivalence
$\iota_{*}:\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$
and hence in particular $c_{\iota,\mathrm{Inj}}\cong\iota_{*}$ is an isomorphism. ∎
@@
$c_{\iota,\mathrm{Inj}}:\mathrm{HH}^{*}\left(\mathrm{Qcoh}\left(X\right),\mathcal{M}\right)\xrightarrow{\sim}\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathrm{Inj},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)\right)\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
@@
Corollary 3.8.
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives, $\mathcal{M}$ a $𝕜\Bbbk$-central $\mathcal{X}$-bimodule and let $0\neq\eta\in\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)$. Then we have that its image under the $\mathrm{Inj}$-equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}\left(\eta\right)$ does not vanish.
@@ 4. Construction of $\psi_{\eta}$
We recall the construction of the candidate non-Fourier-Mukai functor $\psi_{\eta}$. In particular from now on we will exclusively consider a functor of the shape $\psi_{\eta}$ for a smooth divisor $X\hookrightarrow Y$ with $\dim X=n$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $M$ a line-bundle, $m\geq n+3$ and $f_{*}\eta=0$.
@@
[LVdB11, Definition 4.2] Let $X=\bigcup_{i=0}^{k}U_{i}$ be an open affine cover of a quasiprojective scheme $X$. Consider for $I\subset\left\{1,...,k\right\}$ the open sets $U_{I}:=\bigcap_{i\in I}U_{i}$. Then $\mathcal{X}$ is the category with objects $\mathrm{P}\left\{1,...,k\right\}\setminus\emptyset$, morphisms:
@@
and composition of morphisms induced by multiplication in the local ring composed with the restriction morphism.
@@
[Kü22, Lemma 5.2] Let $X$ be smooth quasi-projective. Then there is a fully faithful embedding
@@
[Kü22, Lemma 5.4] Let $f:X\hookrightarrow Y$ be a closed embedding of quasi projective schemes. Then we may pull back the open affine cover of $Y$ in the construction of $\mathcal{Y}$ and get an induced functor $\mathfrak{f}:\mathcal{Y}\to\mathcal{X}$ such that the diagram
@@
As we will mostly be working with the category $\mathcal{D}^{b}\left(\mathcal{X}\right)$ and $w$ is an exact embedding we will denote the functor $\mathfrak{f}:\mathcal{X}\to\mathcal{Y}$ by $f:\mathcal{X}\to\mathcal{Y}$ for ease of readability. We also use the slightly strange convention of $f_{*}$ for the pullback on the $𝕜\Bbbk$-linear side in order to mimic the geometric setting and improve readability.
@@
[Kü22, Proposition 5.5] Let $X$ be smooth projective of dimension $n$ and let $\eta\in\mathrm{HH}^{\geq n+3}\left(X,M\right)$. Then there exists an exact functor
@@
Construction 4.6 works out as well if $\eta$ is zero, however, the arising functor then trivially admits a lift to the $\mathcal{A}_{\infty}$-level and so we will not consider this case.
@@ 5. Recollection on Obstructions
Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category, assume $\mathcal{M}$ is right-flat, $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ and let $U\in\mathcal{X}\mathrm{-mod}$.
@@
Dually, if $\mathcal{M}$ is left-projective and $U\in\mathcal{X}\mathrm{-mod}$. A colift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
@@
Let $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$ for a small $𝕜\Bbbk$-linear category $\mathcal{I}$. Let $\widehat{\iota}\in\mathcal{D}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\right)$. Then the following are equivalent:
  • The object $\widehat{\iota}$ is a colift of $\iota\in\mathcal{D}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$.
  • The object $\widehat{\iota}$ is an $\mathcal{A}_{\infty}$-lift of $L\circ\iota\in\mathcal{D}\left(\mathcal{X}_{\eta}\right)_{\mathcal{I}}$ i.e. $L\circ\iota:\mathcal{I}\to\mathcal{X}_{\eta}$.
@@
$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)$ $\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op},\iota\right)$
@@
In particular such a colift cannot exist, and hence no $\mathcal{A}_{\infty}$-lift as well. ∎
@@ 6. Structure of the functor $\widetilde{f}$
It will be key for our conclusion that we carefully pick the space we consider for computing our obstructions, respectively understand the structure of the functor $\widetilde{f}$. In particular we will from now on fix the inclusion of a smooth divisor $f:X\hookrightarrow Y$, a pulled back linebundle $M=f^{*}\widehat{M}$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $m\geq\dim X+3$.
@@
Consider the distinguished triangle in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ from [RVdBN19, Lemma 11.3]
@@
$\mathrm{Ext}^{1}_{\mathcal{Y}}\left(\Sigma^{2-m}f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\cong\mathrm{Ext}^{m-1}_{\mathcal{Y}}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$ $\mathrm{Ext}^{i}\left(\Sigma^{-j}\_,\_\right)\cong\mathrm{Ext}^{i+j}\left(\_,\_\right)$
@@
$\mathrm{Ext}_{Y}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ $\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\left(f_{*}I\otimes f_{*}M\left[2-m\right]\right),f_{*}J\oplus f_{*}\left(J\otimes M\left[2-m\right]\right)\right)$
$\cong\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}\left(J\otimes M\right)[2-m]\right)\\ \mathrm{Ext}_{Y}^{*}\left(f_{*}\left(I\otimes M\right)[2-m],f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\end{pmatrix}$
@@
$Ext_{Y}^{*}\left(f_{*}F,f_{*}G\right)$ $\cong\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}F,G\right)$
$\cong\mathrm{Ext}^{*}_{X}\left(I\otimes C_{f}[-1]\oplus F,G\right)$ $f^{*}f_{*}F\cong\left(F\oplus F\otimes C_{f}[1]\right)$
$\cong\mathrm{Ext}^{*}_{X}\left(F,G\right)\oplus\mathrm{Ext}^{*}_{X}\left(F\otimes C_{f},G\right)$ $\mathrm{Ext}^{*}\text{ is additive}$
@@
In order to have control over the obstructions explicitly, we will construct the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ from Lemma 4.6 explicitly. For this consider the commutative diagram of $\mathcal{A}_{\infty}$-categories and functors giving rise to a good candidate.
@@
The bottom arrow is the functor given in Lemma 4.4 and we denoted by can the canonical functors arising from the infinitesimal deformation.
@@
    $\begin{pmatrix}f&0\\ 0&f\end{pmatrix}:\mathcal{Y}_{f_{*}\eta}\left(i,j\right)=\mathcal{Y}\left(i,j\right)\oplus f_{*}\mathcal{M}\left(i,j\right)[2-m]\to\mathcal{X}\left(i,j\right)\oplus\mathcal{M}\left(i,j\right)[2-m].$
@@
We then use the explicit structure of the category $f_{*}\mathrm{Inj}_{f_{*}\eta}$ to compute that these obstructions cannot vanish independent of intermediate lifts.
Conceptually one can think about this argument as the bimodule $\psi_{\eta}$ not being $\mathcal{A}_{\infty}$ over $\mathrm{Inj}$, however, it is $\mathcal{A}_{\infty}$ over $\mathcal{Y}$. In particular it behaves like a non-$\mathcal{A}_{\infty}$ Fourier-Mukai kernel
The following Propositions 6.5 and Proposition 6.6 shows that the notations $\mathrm{Inj}_{\eta}$ and $f_{*}\mathrm{Inj}_{f_{*}\eta}$ for the essential image of $f_{*}\circ L\circ\iota$ are warranted.
@@
By construction of $L$, in particular [RVdBN19, (5.7)], the morphism spaces in $\mathrm{Inj}_{\eta}$ are of the shape
@@
The essential image of $f_{*}\circ L\circ\iota$ has objects $i\in I$ and morphism spaces given by
@@
As $h$ is a homotopy equivalence the same holds for $\mathrm{essim}\left(h_{*}\circ f_{*}\circ L\circ\iota\right)$. By Lemma 6.1 we have $\psi_{\eta}\left(I\right)\cong f_{*}I\oplus f_{*}\left(M\otimes I\right)[2-m]$. As we made the choice for $s$ being the trivial inclusion $\mathcal{Y}\hookrightarrow\mathcal{Y}_{0}\cong\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$, $s_{*}$ induces the inclusions of upper triangular matrices with identical diagonal, as claimed. ∎

### v2

@@ 3.2. The $\mathrm{Inj}$-equivariant characteristic morphism
$\iota:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}\left(\mathcal{C}\right),$
for $\mathcal{C}$ a category with enough injectives.
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the inclusion $\iota_{*}:\mathrm{Inj}\hookrightarrow\mathcal{C}\subset\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$ is dg.
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the functor $\iota^{*}:\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$ is a quasi-equivalence.
@@
Let $\mathcal{C}$ be a $𝕜\Bbbk$-linear category with enough injectives and finite global dimension. Then the equivariant characteristic morphism
$c_{\iota,\mathrm{Inj}}\cong\iota_{*}:\mathrm{HH}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\mathrm{Inj}}}\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
is an isomorphism.
@@
By Section 3.1 above we can interpret the equivariant characteristic morphism as the pullback along the inclusion of the full subcategory
$\iota:\mathrm{Inj}\to\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)$
in the homotopy category of $\mathrm{dg}$-categories. In particular we need to pick a fibrant replacement of $\mathrm{Inj}$. By [Tab07, Lemme 2.13] pretriangulated replacements are an incarnation of a fibrant replacement. So we may choose the canonical pretriangulated hull given by $\mathrm{dg}$-modules $\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Inj}\right)$. By Proposition 3.4 $\iota$ induces an equivalence
and hence in particular
$c_{\iota,\mathrm{Inj}}\cong\iota_{*}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}^{*}_{\mathcal{C}_{\mathrm{Inj}}}\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
is an isomorphism. ∎
@@
$c_{\iota,\mathrm{Inj}}:\mathrm{HH}^{*}\left(\mathrm{Qcoh}\left(X\right),\mathcal{M}\right)\xrightarrow{\sim}\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathrm{Inj},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathrm{Qcoh}\left(X\right)\right)\right)\left(\iota,\psi_{\mathcal{M}}\circ\iota\right)$
@@ 4. Construction of $\psi_{\eta}$
We recall the construction of the candidate non-Fourier-Mukai functor $\psi_{\eta}$.
@@
[LVdB11, Definition 4.2] Let $X=\bigcup_{i=0}^{k}U_{i}$ be an open affine cover of a quasiprojective scheme $X$. Consider for $\emptyset\neq I\subset\left\{1,...,k\right\}$ the open sets $U_{I}:=\bigcap_{i\in I}U_{i}$. Then $\mathcal{X}$ is the category with objects $\mathrm{P}\left\{1,...,k\right\}\setminus\emptyset$, morphisms:
@@
and composition of morphisms induced by multiplication in the local rings composed with the restriction morphism.
@@
[Kü24, Lemma 5.2] Let $X$ be smooth quasi-projective. Then there is a fully faithful embedding
@@
Corollary 4.4.
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category associated to a scheme, let $\mathcal{M}=WM$ for $M$ a quasicoherent sheaf and let $0\neq\eta\in\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)$. Then we have that its image under the $\mathrm{Inj}$-equivariant characteristic morphism $c_{\iota,\mathrm{Inj}}\left(\eta\right)$ does not vanish.
@@
[Kü24, Lemma 5.4] Let $f:X\hookrightarrow Y$ be a closed embedding of quasi projective schemes. Then we may pull back the open affine cover of $Y$ in the construction of $\mathcal{Y}$ and get an induced functor $\mathfrak{f}:\mathcal{Y}\to\mathcal{X}$ such that the diagram
@@
As we will mostly be working with the category $\mathcal{D}^{b}\left(\mathcal{X}\right)$ and $w$ is an exact embedding we will denote the functor $\mathfrak{f}:\mathcal{X}\to\mathcal{Y}$ by $f:\mathcal{X}\to\mathcal{Y}$ for ease of readability.
@@
[Kü24, Proposition 5.5] Let $X$ be smooth projective of dimension $n$ and let $\eta\in\mathrm{HH}^{\geq n+3}\left(X,M\right)$. Then there exists an exact functor
@@
Construction 4.7 works as well if $\eta$ is zero, however, the arising functor then trivially admits a lift to the $\mathcal{A}_{\infty}$-level and so we will not consider this case.
@@ 5. Recollection on Obstructions
Let $\mathcal{X}$ be a small $𝕜\Bbbk$-linear category, assume that $\mathcal{M}$ is a right-flat $𝕜\Bbbk$-central bimodule, $\eta\in\mathrm{HH}^{\geq 3}\left(\mathcal{X},\mathcal{M}\right)$ and let $U\in\mathcal{X}\mathrm{-mod}$.
@@
Dually, if $\mathcal{M}$ is a left-projective $𝕜\Bbbk$-central bimodule and $U\in\mathcal{X}\mathrm{-mod}$. A colift of $U$ to $\mathcal{X}_{\eta}$ is an object $V\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\right)$, such that there exists an isomorphism in $\mathcal{X}_{\eta}\mathrm{-mod}$:
@@
Let $\iota\in\mathcal{D}^{b}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$ for a small $𝕜\Bbbk$-linear category $\mathcal{I}$. Let $\widehat{\iota}\in\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op}\right)$. Then the following are equivalent:
  • The object $\widehat{\iota}$ is a colift of $\iota\in\mathcal{D}^{b}\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)$.
  • The object $\widehat{\iota}$ is an $\mathcal{A}_{\infty}$-lift of $L\circ\iota\in\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)_{\mathcal{I}}$ i.e. $L\circ\iota:\mathcal{I}\to\mathcal{X}_{\eta}$.
@@
$\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\left(\mathcal{X}\otimes\mathcal{I}^{op}\right)_{\eta\cup 1},\iota\right)$
$\cong\mathrm{Ext}^{*}_{\mathcal{X}\otimes\mathcal{I}^{op}}\left(\mathcal{X}_{\eta}\otimes\mathcal{I}^{op},\iota\right)$
@@
In particular such a colift cannot exist, and hence there also cannot exist an $\mathcal{A}_{\infty}$-lift. ∎
@@ 6. Structure of the functor $\widetilde{f}$
It will be key for our conclusion that we carefully pick the space we consider for computing our obstructions, respectively understand the structure of the functor $\widetilde{f}$. In particular we will from now on fix the inclusion of a smooth divisor $f:X\hookrightarrow Y$, a pulled back linebundle $M=f^{*}\widehat{M}$ and $0\neq\eta\in\mathrm{HH}^{m}\left(X,M\right)$ with $m\geq\dim X+3$ such that $f_{*}\eta$.
@@
Consider the distinguished triangle in $\mathcal{D}^{b}\left(\mathcal{X}_{\eta}\right)$ induced by $L$, see [RVdBN19, Lemma 11.3]
@@
$\mathrm{Ext}^{1}_{\mathcal{Y}}\left(\Sigma^{2-m}f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$
$\cong\mathrm{Ext}^{m-1}_{\mathcal{Y}}\left(f_{*}\mathcal{F}\otimes\mathcal{M},f_{*}\mathcal{F}\right)$
@@
$\mathrm{Ext}_{Y}^{*}\left(\psi_{\eta}I,\psi_{\eta}J\right)$ $\cong\mathrm{Ext}_{Y}^{*}\left(f_{*}I\oplus\left(f_{*}I\otimes\widehat{M}\right)\left[2-m\right],f_{*}J\oplus\left(f_{*}J\otimes\widehat{M}\right)\left[2-m\right]\right)$
$\cong\begin{pmatrix}\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}\left(J\right)\otimes\widehat{M}[2-m]\right)\\ \mathrm{Ext}_{Y}^{*}\left(f_{*}\left(I\right)\otimes\widehat{M}[2-m],f_{*}J\right)&\mathrm{Ext}_{Y}^{*}\left(f_{*}I,f_{*}J\right)\end{pmatrix}$
@@
$\mathrm{Ext}_{Y}^{*}\left(f_{*}F,f_{*}G\right)$ $\cong\mathrm{Ext}^{*}_{X}\left(f^{*}f_{*}F,G\right)$
$\cong\mathrm{Ext}^{*}_{X}\left(I\otimes C_{f}[-1]\oplus F,G\right)$ $\mathllap{f^{*}f_{*}F\cong\left(F\oplus F\otimes C_{f}[1]\right)}$
$\cong\mathrm{Ext}^{*}_{X}\left(F,G\right)\oplus\mathrm{Ext}^{*}_{X}\left(F\otimes C_{f},G\right)$ $\quad\quad\mathrm{Ext}^{*}\text{ is additive}$
@@
In order to have control over the obstructions explicitly, we will study the functor $\widetilde{f}:\mathcal{Y}\to\mathcal{X}_{\eta}$ from Lemma 4.7 explicitly. For this consider the commutative diagram of $\mathcal{A}_{\infty}$-categories and functors giving rise to a candidate.
@@
The bottom arrow is the functor given in Lemma 4.5 and we denote by can the canonical functors arising from the infinitesimal deformation, see Definition 2.5.
@@
    $\begin{pmatrix}f&0\\ 0&f\end{pmatrix}:\mathcal{Y}\left(i,j\right)\oplus f_{*}\mathcal{M}\left(i,j\right)[2-m]\to\mathcal{X}\left(i,j\right)\oplus\mathcal{M}\left(i,j\right)[2-m].$
@@
Conceptually one can think about this argument as the bimodule $\psi_{\eta}$ not being $\mathcal{A}_{\infty}$ over $\mathrm{Inj}$ while being $\mathcal{A}_{\infty}$ over $\mathcal{Y}$. In particular it behaves like a non-$\mathcal{A}_{\infty}$ Fourier-Mukai kernel.
The following Proposition 6.5 and Proposition 6.6 show that the notations $\mathrm{Inj}_{\eta}$ and $f_{*}\mathrm{Inj}_{f_{*}\eta}$ for the essential image of $f_{*}\circ L\circ\iota$ are warranted.
@@
By construction of $L$, see [RVdBN19, (5.7)], the morphism spaces in $\mathrm{Inj}_{\eta}$ are of the shape
@@
The essential image of $f_{*}\circ L\circ\iota$ has objects $I\in\mathrm{Inj}$ and morphism spaces given by
@@
As $h$ is a homotopy equivalence the same holds for $\mathrm{essim}\left(h_{*}\circ f_{*}\circ L\circ\iota\right)$. By Lemma 6.1 we have $\psi_{\eta}\left(I\right)\cong f_{*}I\oplus f_{*}\left(M\otimes I\right)[2-m]$. As we made the choice for $s$ being the trivial inclusion $\mathcal{Y}\hookrightarrow\mathcal{Y}_{0}\cong\mathcal{Y}\oplus f_{*}\mathcal{M}[2-m]$, $s_{*}$ induces the inclusions of upper triangular matrices with identical diagonal as claimed. ∎